OpenStudy (anonymous):

the number of cells in a tumor doubles every 6 weeks starting with 8 cells. after how many weeks does the tumor have 1500 cells?

OpenStudy (anonymous):

this is a geometric series with common ratio 2 sum = 1500 = 8 2^n - 1 / 2-1 1501 = 8* 2^n - 8 2^n = 1501 / 8 = n log 2 = log 187.6 n = 7.55 number weeks = 7.55 * 6 = 45.3 weeks

OpenStudy (anonymous):

3rd line should be 1501 = 8* 2^n

OpenStudy (anonymous):

sum of a G S = a x r^n - 1 ------- r - 1 where a = first term and r = common ratio

OpenStudy (anonymous):

r u ok with this jessica? any questions?

OpenStudy (anonymous):

Yes, thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

plz click good answer to close question